# Floquet-Engineering Weyl Points and Linked Fermi Arcs from Straight Nodal Lines

### v2

@@ Abstract
Floquet engineering provides a powerful and flexible method for modifying the band structures of quantum materials. While circularly polarized light has been shown to convert curved nodal lines in three-dimensional semimetals into Weyl points, such a transformation is forbidden for an isolated straight nodal line. In this work, we uncover a dramatic shift in this paradigm when multiple straight nodal lines intersect. We observe that circularly polarized light not only gaps them into Weyl points but also induces unprecedented surface-state Fermi arcs that extend across the entire surface Brillouin zone and form a linked topological structure. These findings advance our fundamental understanding of light-driven transitions in topological semimetals and unveil a unique Weyl semimetal phase defined by linked Fermi arcs. We discuss potential exotic phenomena arising from this phase, applications of our predictions to spin-splitting antiferromagnets, and the extension of this Weyl semimetal phase to classical systems.
@@
Introduction.—A groundbreaking advance in Floquet engineering emerged with the discovery that circularly polarized light (CPL) can open a band gap at graphene’s Dirac points [1], realizing topological gapped phases now known as Floquet Chern insulators. This seminal work ignited widespread investigation into Floquet topological phases [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] and light-driven topological phase transitions in quantum materials [12, 13, 14, 15]. The direct observation of light-driven band structure modifications, however, was challenging for years until recent notable advances in time- and angle-resolved photoemission spectroscopy (TrARPES). This technique’s unprecedented capability to resolve photon-dressed band structure in momentum space [16, 17, 18, 19] is now propelling the field into an exciting new phase of discovery.
@@
Band degeneracy serves as the source for the nontrivial topological properties of the band structure [20]. The influence of CPL on band degeneracy depends critically on both the system’s dimension and the structure of the degeneracy. In two-dimensional (2D) systems, the most common type of band degeneracy is Dirac points protected by $\mathcal{PT}$ symmetry. Since CPL generally breaks $\mathcal{PT}$ symmetry, 2D Dirac points are unstable under CPL irradiation. Consequently, the 2D Floquet topological phases induced by CPL typically exhibit gapped band structures [21, 22, 23]. In three-dimensional (3D) systems, band degeneracies exhibit greater diversity, encompassing nodal points [24, 25, 26, 27, 28], nodal lines [29, 30, 31, 32, 33, 34, 35, 36, 37], and nodal surfaces [38, 39]. This diversity leads to a richer interplay between CPL and band degeneracy compared to 2D systems. Among these band degeneracies, Weyl points are special because their stability is topologically protected [40], unlike other band degeneracies that rely on symmetry protection. When subjected to CPL, Weyl points primarily experience chirality-dependent positional shifts [41], while other degeneracies typically transform into Weyl points [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. This conversion results in a robust topological gapless phase—the Floquet Weyl semimetal. Particularly interesting is the transition from curved nodal lines to Weyl points [illustrated in Fig. 1(b)], which typically generates widely separated Weyl points [44, 45, 46] and enables the realization of multi-Weyl points [48, 49]. This large separation leads to extended Fermi-arc surface states at the boundary and gives rise to pronounced anomalous Hall effects.
@@
Crystalline symmetries can enforce nodal lines to be straight [62, 63, 64, 65, 66, 67, 68, 69, 70]. However, research on Floquet Weyl points (FWPs) arising from nodal lines has thus far been confined to curved cases [44, 45, 46, 47, 48, 49, 50, 51, 52]. This restriction stems from the fact that a straight nodal line (SNL) effectively represents a stack of 2D Dirac points along a specific momentum direction. Since these Dirac points are identical, CPL affects them uniformly—either gapping all of them simultaneously or leaving the entire stack unaffected (as demonstrated below). Consequently, this uniform response prevents the generation of FWPs from SNLs [illustrated in Fig. 1(c)]. In this work, we demonstrate that this paradigm undergoes a dramatic transformation when multiple SNLs intersect. Beyond the unexpected emergence of FWPs, we discover an extraordinary characteristic of the resulting Fermi-arc surface states: they span the entire surface Brillouin zone (BZ), forming an intricately linked structure—a phenomenon never before observed in either static or Floquet Weyl semimetals. This unprecedented feature originates from a nontrivial pattern in the layer Chern numbers. We note that linked Fermi loops have recently been experimentally realized in photonic 3D Chern insulators [71]. However, a crucial distinction exists between these fully gapped Chern insulators and our Floquet Weyl semimetals. In the former, the layer Chern numbers remain constant along a specific momentum direction, whereas in the latter, they exhibit sign reversal. We also investigate the anomalous Hall effect in this system, uncovering a strong dependence on the propagation direction of the incident light.
@@
where $\mathcal{H}_{n}$ denotes the $n$th-decomposition coefficient matrix, i.e., $\mathcal{H}(\boldsymbol{k},t)=\sum_{n}\mathcal{H}_{n}(\boldsymbol{k})e^{in\omega t}$ with $n\in\mathbb{Z}$. For this Hamiltonian, we have $\mathcal{H}_{0}(\boldsymbol{k})=k_{y}\sigma_{x}+k_{z}\sigma_{y}$ and ±$\mathcal{H}_{\pm 1}(\boldsymbol{k})=-\frac{eA_{0}}{2}[(\cos\phi\pm i\eta\sin\phi\cos\theta)\sigma_{x}\mp i\eta\sin\theta\sigma_{y}]$, which yields:
@@
[…]
@@
The emergence of FWPs reveals that intersected SNLs host fundamentally distinct physical behavior compared to their isolated counterparts.
The FWPs emerging from intersected SNLs exhibit several striking differences compared to those generated in other systems. Specifically, we observe that: (i) A distinctive scaling of Weyl-point separation ($d_{w}$). For FWPs originating from a Dirac point, the separation scales as $d_{w}\propto A_{0}^{2}/\omega$ [42]; For a circular nodal ring, $d_{w}$ is approximately equal to the ring’s diameter [illustrated in Fig. 1(b)] and thus independent of $A_{0}$ and $\omega$ to the leading order [44, 45, 46]; In the present case, the scaling follows $d_{w}\propto A_{0}$, distinguishing it from both Dirac points and nodal rings. (ii) A distinctive dependence of the displacement vector of FWPs on the light’s direction. When the light propagates along $(\phi,\theta)$ on the Bloch sphere, the displacement vector aligns with the light direction for the case of Dirac points [42]. For a circular nodal ring, it aligns with $(\phi,\frac{\pi}{2})$, corresponding to the in-plane projection of the light’s orientation [44, 45, 46], as illustrated in Fig. 1(b). In stark contrast, our analysis [Eq. (7)] reveals that for intersected SNLs, the vector is oriented along $(\frac{\pi}{2}-\phi,\frac{\pi}{2})$, which generally deviates from the light’s orientation or its projection, as illustrated in Fig. 1(d).
@@
The positions of FWPs are simply determined by the zeros of the coefficients in front of the Pauli matrices. Without loss of generality, we focus on a specific case for a detailed analysis. Specifically, we consider light propagating along the direction $(\theta,\phi)=(\frac{\pi}{3},\frac{\pi}{3})$. The resulting positions and chiralities of FWPs are illustrated in Fig. 2(b). As expected, we observe four pairs of FWPs in the $k_{z}=0$ plane and another four pairs in the $k_{z}=\pi$ plane. Notably, the distributions of positions and chiralities in the two $k_{z}$ planes are identical. Consequently, when the bulk Weyl points are projected along the $z$ direction, those with the same topological charge overlap, yielding a net topological charge of $+2$ or $-2$ per Weyl-point projection. Each Weyl-point projection is thus accompanied by two Fermi arcs, a property we will demonstrate below.
@@
The unique momentum dependence of $C(k_{x})$ and $C(k_{y})$ indicates that the Fermi arcs on the $z$-normal surfaces will exhibit a linked configuration. To demonstrate this, we need to determine the spectrum of the surface states. This can be done by numerically calculating the energy spectrum for a sample with periodic boundary conditions in the $x$ and $y$ directions and open boundary conditions in the $z$ direction (see Sec. II of the SM [75]). Notably, for the current Hamiltonian, the surface-state spectra can be analytically obtained as the Floquet Hamiltonian in Eq. (9) takes a form similar to the Su-Schrieffer-Heeger model in the $z$ direction [78]. Specially, the surface-state spectra for the top (+)(+)( + ) and bottom ()(-)( - ) $z$-normal surfaces are given by
@@
where the function $\text{sgn}(x)$ equals $1$ $(-1)$ for $x>0$ ($x<0$). Fig. 3(b) displays the spectrum for the top surface. By tracing the zero-energy contours and Weyl-point projections, we obtain the connectivity of the Fermi arcs, as depicted in Fig. 3(c). Indeed, each Weyl-point projection connects to two Fermi arcs, one from each side. Interestingly these Fermi arcs collectively span the entire surface BZ. When the 2D surface BZ is compactified into a torus by identifying opposite boundaries, the Fermi arcs clearly organize into two mutually linked loops, as illustrated in Fig. 3(d). While Fermi arcs spanning the entire surface BZ have been experimentally observed in chiral topological semimetals [79, 80, 81, 82], the linked structure we find here represents a novel feature that has not been previously identified in either static or Floquet Weyl semimetals.
@@
where $G^{<}_{nm,ij}$ is the Floquet lesser Green’s function and $v^{y}_{mn,ji}(\mathbf{k}_{s})$ is the Floquet velocity matrix elements (see the End Matter for details). The indices $\{n,m\}$ label the sector in the frequency space, while $\{i,j\}$ label lattice sites along the $z$ direction. Repeated indices are implicitly summed over. Once the current is determined, the Hall conductivity $\sigma_{yz}$ is calculated as $\sigma_{yz}=J_{y}/E_{z}=J_{y}N_{z}/V$.
@@
Although intersected SNLs can appear in nonmagnetic and magnetic materials, here we highlight the relevance of our theoretical predictions to spin-splitting antiferromagnets—a class of materials exhibiting momentum-dependent spin splitting and symmetry-enforced zero-net magnetization  [92, 93, 94, 95, 96, 97, 98, 99, 100]. The underlying symmetries responsible for these characteristics naturally stabilize SNLs along high symmetry lines in the BZ [66, 70]. Notably, recent theoretical work has identified several existing compounds—including $\beta$-Fe2(PO4)O, Co2(PO4)O and LiTi2O4—as altermagnets hosting such intersected SNLs [67]. Based on our study, a light-induced topological transition from SNLs to Weyl points can be probed by TrARPES in these materials. Beyond quantum materials, the predicted Weyl phase characterized by linked Fermi arcs also offers exciting opportunities for investigating unconventional bulk-boundary physics in classical systems [101, 102, 103, 104, 105, 106]. The minimal two-band Hamiltonian in Eq. (9) also serves as an ideal model for experimental exploration due to its simplicity, enabling direct implementation in classical systems.

### v3

# Floquet Weyl Semimetals with Linked Fermi Arcs

@@ Abstract
Floquet engineering provides a powerful and flexible method for modifying the band structures of quantum materials. While circularly polarized light has been shown to convert curved nodal lines in three-dimensional semimetals into Weyl points, such a transformation is forbidden for an isolated straight nodal line. In this work, we uncover a dramatic shift in this paradigm when multiple straight nodal lines intersect. We observe that circularly polarized light not only gaps them into Weyl points but also induces unprecedented surface-state Fermi arcs that extend across the entire surface Brillouin zone and form a linked topological structure. These findings advance our fundamental understanding of light-driven transitions in topological semimetals and unveil a unique Weyl semimetal phase defined by linked Fermi arcs. We discuss potential exotic phenomena arising from this phase, applications of our predictions to spin-split antiferromagnets, and the extension of this Weyl semimetal phase to classical systems.
Keywords: Straight nodal line; Floquet Weyl semimetal; Linked Fermi arc; Spin-split antiferromagnet.
@@
Introduction.—A groundbreaking advance in Floquet engineering emerged with the discovery that circularly polarized light (CPL) can open a band gap at graphene’s Dirac points [1], realizing topological gapped phases now known as Floquet Chern insulators. This seminal work ignited widespread investigation into Floquet topological phases [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] and light-driven topological phase transitions in quantum materials [12, 13, 14, 15]. The direct observation of light-driven band structure modifications, however, was challenging for years until recent notable advances in time- and angle-resolved photoemission spectroscopy (TrARPES). This technique’s unprecedented capability to resolve photon-dressed band structure in momentum space [16, 17, 18, 19, 20] is now propelling the field into an exciting new phase of discovery.
@@
Band degeneracy serves as the source for the nontrivial topological properties of the band structure [21]. The influence of CPL on band degeneracy depends critically on both the system’s dimension and the structure of the degeneracy. In two-dimensional (2D) systems, the most common type of band degeneracy is Dirac points protected by $\mathcal{PT}$ symmetry. Since CPL generally breaks $\mathcal{PT}$ symmetry, 2D Dirac points are unstable under CPL irradiation. Consequently, the 2D Floquet topological phases induced by CPL typically exhibit gapped band structures [22, 23, 24]. In three-dimensional (3D) systems, band degeneracies exhibit greater diversity, encompassing nodal points [25, 26, 27, 28, 29], nodal lines [30, 31, 32, 33, 34, 35, 36, 37, 38], and nodal surfaces [39, 40]. This diversity leads to a richer interplay between CPL and band degeneracy compared to 2D systems. Among these band degeneracies, Weyl points are special because their stability is topologically protected [41], unlike other band degeneracies that rely on symmetry protection. When subjected to CPL, Weyl points primarily experience chirality-dependent positional shifts [42, 43], while other degeneracies typically transform into Weyl points [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64]. This conversion results in a robust topological gapless phase—the Floquet Weyl semimetal. Particularly interesting is the transition from curved nodal lines to Weyl points [illustrated in Fig. 1(b)], which typically generates widely separated Weyl points [46, 47, 48] and enables the realization of multi-Weyl points [50, 51]. This large separation leads to extended Fermi-arc surface states at the boundary and gives rise to pronounced anomalous Hall effects.
@@
Crystalline symmetries can enforce nodal lines to be straight [65, 66, 67, 68, 69, 70, 71, 72, 73]. However, research on Floquet Weyl points (FWPs) arising from nodal lines has thus far been confined to curved cases [46, 47, 48, 49, 50, 51, 52, 53, 54]. This restriction stems from the fact that a straight nodal line (SNL) effectively represents a stack of 2D Dirac points along a specific momentum direction. Since these Dirac points are identical, CPL affects them uniformly—either gapping all of them simultaneously or leaving the entire stack unaffected (as demonstrated below). Consequently, this uniform response prevents the generation of FWPs from SNLs [illustrated in Fig. 1(c)]. In this work, we demonstrate that this paradigm undergoes a dramatic transformation when multiple SNLs intersect. Beyond the unexpected emergence of FWPs, we discover an extraordinary characteristic of the resulting Fermi-arc surface states: they span the entire surface Brillouin zone (BZ), forming an intricately linked structure—a phenomenon never before observed in either static or Floquet Weyl semimetals. We also investigate the anomalous Hall effect in this system, uncovering a strong dependence on the propagation direction of the incident light.
@@
where $\mathcal{H}_{n}$ denotes the $n$th-decomposition coefficient matrix, i.e., $\mathcal{H}(\boldsymbol{k},t)=\sum_{n}\mathcal{H}_{n}(\boldsymbol{k})e^{in\omega t}$ with $n\in\mathbb{Z}$. For this Hamiltonian, we have $\mathcal{H}_{0}(\boldsymbol{k})=k_{y}\sigma_{x}+k_{z}\sigma_{y}$ and $\mathcal{H}_{\pm 1}(\boldsymbol{k})=-\frac{eA_{0}}{2}[(\cos\phi\pm i\eta\sin\phi\cos\theta)\sigma_{x}\mp i\eta\sin\theta\sigma_{y}]$, which yields:
@@
Although FWPs cannot be generated, fully gapping out the SNL instead results in a 3D Floquet Chern insulator characterized by layer Chern numbers $(C(k_{x}),C(k_{y}),C(k_{z}))$. Here, $C(k_{i})$ denotes the layer Chern number defined on the 2D momentum plane with fixed $k_{i}$, given explicitly by $C(k_{i})=\frac{1}{2\pi}\sum_{j,l}\epsilon_{ijl}\int dk_{j}dk_{l}\partial_{j}A_{l}$ where $\epsilon_{ijl}$ is the third-order antisymmetry tensor, and $A_{l}=i\langle u(\boldsymbol{k})|\partial_{l}u(\boldsymbol{k})\rangle$ is the Berry connection of the occupied band. For the continuum Hamiltonian considered here, we find
@@
The emergence of FWPs reveals that intersected SNLs host fundamentally distinct physical behavior compared to their isolated counterparts. In Eq. (7), a notable property is that the positions of FWPs depend solely on the amplitude $A_{0}$ and are independent of the frequency $\omega$. This independence implies a remarkable robustness: the high-frequency expansion accurately predicts the Weyl-point positions even when extended into lower-frequency regimes where it fails to capture the full band structure.
The FWPs emerging from intersected SNLs exhibit several striking differences compared to those generated in other systems. Specifically, we observe that: (i) A distinctive scaling of Weyl-point separation ($d_{w}$). For FWPs originating from a Dirac point, the separation scales as $d_{w}\propto A_{0}^{2}/\omega$ [44]. Such a scaling is also applied to other nodal point semimetals [64]. For a circular nodal ring, $d_{w}$ is approximately equal to the ring’s diameter [illustrated in Fig. 1(b)] and thus independent of $A_{0}$ and $\omega$ to the leading order [46, 47, 48]; In the present case, the scaling follows $d_{w}\propto A_{0}$, distinguishing it from both Dirac points and nodal rings. (ii) A distinctive dependence of the displacement vector of FWPs on the light’s direction. When the light propagates along $(\phi,\theta)$ on the Bloch sphere, the displacement vector aligns with the light direction for the case of Dirac points [44]. For a circular nodal ring, it aligns with $(\phi,\frac{\pi}{2})$, corresponding to the in-plane projection of the light’s orientation [46, 47, 48], as illustrated in Fig. 1(b). In stark contrast, our analysis [Eq. (7)] reveals that for intersected SNLs, the vector is oriented along $(\frac{\pi}{2}-\phi,\frac{\pi}{2})$, which generally deviates from the light’s orientation or its projection, as illustrated in Fig. 1(d).
@@
The positions of FWPs are simply determined by the zeros of the coefficients in front of the Pauli matrices. Without loss of generality, we focus on a specific case for a detailed analysis. Specifically, we consider light propagating along the direction $(\theta,\phi)=(\frac{\pi}{3},\frac{\pi}{3})$. The resulting positions and chiralities of FWPs are illustrated in Fig. 2(b). As expected, we observe four pairs of FWPs in the $k_{z}=0$ plane and another four pairs in the $k_{z}=\pi$ plane. Notably, the distributions of positions and chiralities in the two $k_{z}$ planes are identical. Consequently, when the bulk Weyl points are projected along the $z$ direction, those with the same topological charge overlap, yielding a net topological charge of +2+2 or 2-2 per Weyl-point projection. Each Weyl-point projection is thus accompanied by two Fermi arcs, a property we will demonstrate below.
@@
The unique momentum dependence of $C(k_{x})$ and $C(k_{y})$ indicates that the Fermi arcs on the $z$-normal surfaces will exhibit a linked configuration. To demonstrate this, we need to determine the spectrum of the surface states. This can be done by numerically calculating the energy spectrum for a sample with periodic boundary conditions in the $x$ and $y$ directions and open boundary conditions in the $z$ direction (see Sec. II of the SM [77]). Notably, for the current Hamiltonian, the surface-state spectra can be analytically obtained as the Floquet Hamiltonian in Eq. (9) takes a form similar to the Su-Schrieffer-Heeger model in the $z$ direction [81]. Specially, the surface-state spectra for the top $(+)$ and bottom ()(-) $z$-normal surfaces are given by
@@
where the function $\text{sgn}(x)$ equals 11 (1)(-1) for $x>0$ ($x<0$). Fig. 3(b) displays the spectrum for the top surface. By tracing the zero-energy contours and Weyl-point projections, we obtain the connectivity of the Fermi arcs, as depicted in Fig. 3(c). Indeed, each Weyl-point projection connects to two Fermi arcs, one from each side. Interestingly these Fermi arcs collectively span the entire surface BZ. When the 2D surface BZ is compactified into a torus by identifying opposite boundaries, the Fermi arcs clearly organize into two mutually linked loops, as illustrated in Fig. 3(d). While Fermi arcs spanning the entire surface BZ have been experimentally observed in chiral topological semimetals [82, 83, 84, 85], the linked structure we find here represents a novel feature that has not been previously identified in either static or Floquet Weyl semimetals.
@@
where $G^{<}_{nm,ij}$ is the Floquet lesser Green’s function and $v^{y}_{mn,ji}(\mathbf{k}_{s})$ is the Floquet velocity matrix elements (see Sec. IV of the SM [77] for details). The indices $\{n,m\}$ label the sector in the frequency space, while $\{i,j\}$ label lattice sites along the $z$ direction. Repeated indices are implicitly summed over. Once the current is determined, the Hall conductivity $\sigma_{yz}$ is calculated as $\sigma_{yz}=J_{y}/E_{z}=J_{y}N_{z}/V$.
@@
Although Weyl points are robust against perturbations and long Fermi arcs can appear on general surfaces, observing the arcs’ linking structure requires careful sample surface preparation. In our model, the linking structure observed on $z$-normal surfaces arises from two closely-related properties of the Hamiltonian: the overlap of projections of bulk Weyl points with identical topological charges, and layer Chern numbers $|C(k_{x})|$ and $|C(k_{y})|$ that are identically 2 across all gapped planes. However, this linking structure is not robust against the variation of surface orientation. For a general surface orientation, the Weyl-point projections no longer overlap, and the layer Chern numbers vary across the Brillouin zone, causing the linking to disappear. If the cleavage surface only slightly deviates from the $z$-normal direction, the Fermi arcs remain proximate but unlinked. Under such conditions, the essential physics of multiple Weyl orbitals persists. In particular, the effect known as magnetic breakdown can still enable electrons to tunnel between adjacent Weyl-orbital trajectories [95].
Intersected SNLs can appear in both nonmagnetic and magnetic materials. Here we highlight the relevance of our theoretical predictions to spin-split antiferromagnets—a class of materials exhibiting momentum-dependent spin splitting and symmetry-enforced zero-net magnetization  [96, 97, 98, 99, 100, 101, 102, 103, 104]. The underlying symmetries responsible for these characteristics naturally stabilize SNLs along high symmetry lines in the BZ [69, 73]. Notably, recent theoretical work has identified several existing compounds—including $\beta$-Fe2(PO4)O, Co2(PO4)O and LiTi2O4—as altermagnets hosting such intersected SNLs [70]. Based on our study, a light-induced topological transition from SNLs to Weyl points can be probed by TrARPES in these materials. We now provide an estimation to the distance of two neighboring FWPs. Specifically, we consider a mid-infrared laser with frequency $\hbar\omega=280$ meV (here the reduced Planck constant is restored) and amplitude $\mathcal{E}=\omega A_{0}=2.1\times 10^{8}$ V/m [105]. Assuming that the laser is incident along the direction $(\theta,\phi)=(\pi/2,\pi/4)$, we obtain $d_{\omega}=\sqrt{2}eA_{0}/\hbar=\sqrt{2}e\mathcal{E}/\hbar\omega\simeq 1\times 10^{9}$ m-1. For the material candidate $\beta$-Fe2(PO4)O (in-plane lattice constant $a=5.419\times 10^{-10}$ m) [70], this distance corresponds to roughly a tenth of the Brillouin-zone size ($G=2\pi/a=1.2\times 10^{10}$ m-1). This light-induced separation is several times larger than that observed in TaAs [106, 107], confirming its experimental viability.
Beyond quantum materials, the predicted Weyl phase characterized by linked Fermi arcs also offers exciting opportunities for investigating unconventional bulk-boundary physics in classical systems [108, 109, 110, 111, 112, 113]. The minimal two-band Hamiltonian in Eq. (9) also serves as an ideal model for experimental exploration due to its simplicity, enabling direct implementation in classical systems.